# Small-Time Local Controllability of the multi-input bilinear Schrödinger equation thanks to a quadratic term

### v3

@@ 1.6 Small-time continuously approximately reachable vector
Finally, let $(\mathscr{F}_{T})_{T>0}$ be a family of functions from $X\times E_{T}$ to $X$ for $T>0$. The goal is to prove that the map $\mathscr{F}_{T}$ in locally onto. To do this, we use the following definition; this is an equivalent of the notion of tangent vector, but in infinite dimension.
@@
A vector $\xi\in X$ is called a small-time continuously approximately reachable vector if there exists a continuous map $\Xi:[0,+\infty[\to X$ with $\Xi(0)=\xi$ such that for all $T>0$, there exist $C,\rho,s>0$ and a continuous map $z\in(-\rho,\rho)\mapsto u_{z}\in E_{T}$ such that,
@@ 1.7 Plan of the paper
The paper is organized as follow: in section $2$, we present a proof of a theorem of STLC thanks to quadratic terms for control-affine system in finite dimension, as a toy-model for the bilinear Schrödinger equation. In section $3$, we present classical properties about the bilinear Schrödinger equation. Finally, in the section $4$, we give the proof of the main theorem of this article. Some elements of proof are developed in the appendix.
@@ 2.1 Framework and notations
For each $T>T_{1}>0$, $u,v\in L^{1}((T_{1},T),\mathbb{R})$, there exists a unique maximal mild solution to (2.1) with initial data $p$ at time $T_{1}$, which we will denote by $x(\cdot;(u,v),p)$. We will sometimes note $x(\cdot;(u,v),p,T_{1})$. As we are interested in small-time, and small controls, the solution is well-defined up to time $T$. The following concepts were introduced by Beauchard and Marbach in [10], in 2018.
@@
Let $f_{0},f_{1},f_{2}$ be $\mathcal{C}^{\infty}(\Omega,\mathbb{R}^{d})$ vector fields on an open subset $\Omega$ of $\mathbb{R}^{d}$ and $f=\{f_{0},f_{1},f_{2}\}$.
@@ 2.3 A quadratic sufficient condition for STLC
Let $d,L\geqslant 1$, $f_{0},f_{1},f_{2}$ be analytic vector fields on a neighborhood of 00 in $\mathbb{R}^{d}$, such that $f_{0}(0)=0$. Assume that $r:=\dim\left(S_{1}(f)(0)\right)<d$ and that there exist brackets $b_{r+1},\cdots,b_{d}\in\mathcal{B}_{2,good}$ such that $\underset{i\in\llbracket r+1,d\rrbracket}{\max}|b_{i}|=L$ and
@@ 2.4 Some tools for the proof
Let $\mathcal{B}\subset\operatorname{Br}(X)$ be a Hall set. The coordinates of the second kind associated with $\mathcal{B}$ is the unique family $(\xi_{b})_{b\in\mathcal{B}}$ of functionals +×$\mathbb{R}_{+}\times L^{1}_{\mathrm{loc}}(\mathbb{R}_{+};\mathbb{R})^{2}\to\mathbb{R}$ defined by induction as: for every $t>0$ and $u,v\in L^{1}((0,t),\mathbb{R})$,
  • $\xi_{X_{0}}(t):=t$, $\xi_{X_{1}}(t,u):=\displaystyle\int_{0}^{t}u(s)\mathrm{d}s=u_{1}(t)$ and $\xi_{X_{2}}(t,v):=\displaystyle\int_{0}^{t}v(s)\mathrm{d}s=v_{1}(t)$,
@@
For every $M\in\mathbb{N}^{*}$, there exists $C_{M}>0$ such that, for every $T\geqslant 0$, $u,v\in L^{1}((0,T),\mathbb{R})$, $b\in\mathcal{B}$ with $n(b)\leqslant M$ and $t\in[0,T]$,
@@ 2.5 A quadratic moment problem
One may assume that $B=\left\{M_{\nu}^{1},\ M_{\nu}^{2},\ C_{\mu,\nu},\ (\mu,\nu)\in\llbracket 0,\mu^{*}\rrbracket\times\llbracket 0,\nu^{*}\rrbracket\right\}.$ One considers $N=\mu^{*}+\nu^{*}+1$. We are looking for $u$ and $v$ in the form of $u=\phi^{(N)}$ and $v=\psi^{(N)}$, with $\phi,\psi\in\mathcal{C}^{\infty}_{c}((0,1),\mathbb{R})$. Then, for every $\mu\in\llbracket 0,\mu^{*}\rrbracket$, $\xi_{M_{\mu}^{1}}(1,(u,v))=u_{\mu+1}(1)=0$. One recalls that for $b\in\mathcal{B}_{1}$, $\eta_{b}=\xi_{b}$. Thus, (2.18) is already verified for $b\in B\cap\mathcal{B}_{1}$.
The equation (2.14) gives: for all $b\in\mathcal{B}_{2,good}\cap B$, $\eta_{b}(1,(u,v))=\xi_{b}(1,(u,v))$. Moreover, using (2.12), one obtains for all $(\mu,\nu)\in\llbracket 0,\mu^{*}\rrbracket\times\llbracket 0,\nu^{*}\rrbracket$,
$\eta_{C_{\mu,\nu}}(1,(u,v))=\int_{0}^{1}\frac{(1-t)^{\nu}}{\nu!}u_{\lfloor\frac{\mu}{2}\rfloor+1}(t)v_{\lfloor\frac{\mu+1}{2}\rfloor}(t)\mathrm{d}t.$
@@
Step 2: Construction of $u$ and $v$. One chooses $u=\phi^{(N)}$, with the function $\phi$ obtained at step $1$. By construction of $\phi$, there exists $\psi\in\mathcal{C}^{\infty}_{c}((0,1),\mathbb{R})$, solution to the following moment problems
@@ 2.6 Tangent vector
Let $m\geqslant 0$, we prove that the system is $W^{m,\infty}_{0}-$STLC. Taking into account the assumption (2.7), we can consider $\mathbb{P}$, the linear projection on $\text{Span}(f_{b_{i}}(0))_{r+1\leqslant i\leqslant d}$ parallel to ${S}_{1}(f)(0)$. The purpose of this section is to prove:
@@
Consequently, the definition of $(\bar{u},\bar{v})$ and (2.24) gives: for all $z\in(-\rho,\rho)$, for all $i\in\{1,2\}$, for all $k\in\llbracket 0,q_{j}\rrbracket$, $\eta_{M_{k}^{i}}(T_{1},(u_{z},v_{z}))=0$ and,
@@
Using the size of the controls given by (2.21) with $(k,r)=(0,1)$ and (2.25), one has:
@@
Let $j\in\llbracket r+1,d\rrbracket$. For all $T_{1}+1>T>T_{1}>0$, there exist $C,\rho>0$ and a continuous map $z\mapsto(U_{z},V_{z})\in W^{m,\infty}_{0}(0,T)^{2}$ such that
@@
Note that $\mathbb{P}_{{S}_{1}(f)(0)}=I-\mathbb{P}$. Then,
@@
$e^{(T-T_{1})H_{0}}\left(I-\mathbb{P}\right)\left(x(T_{1};(u_{z},v_{z}),0,0)\right)\in{S}_{1}(f)(0)=\ker(\mathbb{P}),$
@@
$\begin{gathered}\left\|\mathbb{P}\left(\mathcal{Z}_{2}(T;f,(\tilde{u}_{z},\tilde{v}_{z}))\right)\right\|_{\infty}=\mathcal{O}\left(\left\|(\tilde{u}_{z},\tilde{v}_{z})\right\|_{L^{1}}^{2}\right)=\mathcal{O}\left(|z|^{2\delta_{j}}\right),\\ \left\|\mathcal{Z}_{2}(T;f,(\tilde{u}_{z},\tilde{v}_{z}))\right\|_{\infty},\left\|D\mathcal{Z}_{2}(T;f,(\tilde{u}_{z},\tilde{v}_{z}))\right\|_{\infty}=\mathcal{O}\left(\left\|(\tilde{u}_{z},\tilde{v}_{z})\right\|_{L^{1}}\right)=\mathcal{O}\left(|z|^{\delta_{j}}\right).\end{gathered}$
Therefore, thanks to the estimates (2.38), one has:
@@
as $1+s_{j}\leqslant 2\delta_{j}\leqslant 3\alpha_{j}+3s_{j}$. To conclude, the equations (2.36), (2.37) and (2.39) lead to the (2.30). ∎
@@ 5. $(A_{5})$
    Finally, Proposition 2.24 shows that, for all $j\in\llbracket r+1,d\rrbracket$, $f_{b_{j}}(0)$ is a small-time continuously approximately reachable vector. Then, as a supplementary space of $H$ is given by $\text{Span}\left(\left(f_{b_{j}}(0)\right)_{r+1\leqslant j\leqslant d}\right)$, this condition is verified.
@@ 3 Preliminaries on the Schrödinger equation
In all the document, we will note $\omega_{j}:=\lambda_{j}-\lambda_{1}$, and $\nu_{j}=\lambda_{K}-\lambda_{j}$, with $j\geqslant 1$, $K\geqslant 2$. For $\varphi\in L^{2}(0,1)$, $t\in\mathbb{R}$, one defines
@@ 3.1 Well-posedness
For all $\varphi\in H^{2k}_{(0)}(0,1)$, for all $l\in\{0,\cdots,k-1\}$, one has, thanks to Leibniz formula
@@ 3.2 Expansion of the solution
$\left\{\begin{array}[]{lr}i\partial_{t}\xi=-\partial_{x}^{2}\xi-(u\mu_{1}+v\mu_{2})\Psi,&(0,T)\times(0,1),\\ \xi(\cdot,0)=\xi(t,1)=0,&(0,T),\\ \xi(0,\cdot)=0,&(0,1).\end{array}\right.$ (3.10)
@@
Third step: cubic remainder. We conclude with the same arguments. ∎
@@ 3.3 Control in projection
Let $p,m,k\geqslant 0$, $K\geqslant 2$, such that $k\leqslant p$, $\mu_{1},\mu_{2}$ functions satisfying $\mathbf{(H)_{reg}}$, and $\mathbf{(H)_{lin,K,2}}$. Then, the bilinear Schrödinger equation (1.1) is $H^{m}_{0}-$STLC in projection on $\mathcal{H}$ around the ground state in $H^{2(p+m)+3}_{(0)}$. More precisely, for any $T_{1}<T$, there exists $\delta>0$, $C>0$, and a $\mathcal{C}^{1}$ map $\Gamma_{T_{1},T}:\mathcal{V}_{T_{1}}\times\mathcal{V}_{T}\to H^{m}_{0}((T_{1},T),\mathbb{R})^{2}$, where
@@ 4 Proof of the main theorem
The proof of the main theorem is divided in the following steps:
@@
In order to extract the leading terms of the dynamic of the system (as in finite dimension, with the Magnus-type representation formula), we manipulate the expression given by (3.14). This is the purpose of the following subsection.
@@ 4.1 Asymptotic estimates on the quadratic term of the solution
We define $q:=\lfloor\frac{n}{2}\rfloor$. As a remainder, the sequence $(c_{j})_{j\geqslant 1}$ is defined in Section 1.3 as $c_{j}=\left\langle\mu_{1}\varphi_{1},\varphi_{j}\right\rangle\left\langle\mu_{1}\varphi_{j},\varphi_{K}\right\rangle$. The goal of this subsection is to prove the following proposition.
@@
Let $u,v\in L^{2}((0,T),\mathbb{R})$ be such that, $u_{i}(T)=0$ for $11\leqslant i\leqslant q+$ 1, and $v_{i}(T)=0$ for $11\leqslant i\leqslant q+$ 1. Furthermore, assume that $\mathbf{(H)_{reg}}$, $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$ and $\mathbf{(H)_{quad,K,3}}$ hold. Then,
@@
$R_{T,1}^{q+1}(u):=\displaystyle(-1)^{q}\sum_{j=1}^{+\infty}c_{j}\nu_{j}^{q+1}\omega_{j}^{q+1}I^{j}_{T}(u_{q+1},u_{q+1}),$
$R_{T,2}^{q+1}(v):=\displaystyle(-1)^{q}\sum_{j=1}^{+\infty}\tilde{c_{j}}\nu_{j}^{q+1}\omega_{j}^{q+1}I^{j}_{T}(v_{q+1},v_{q+1}),$
$\rho_{T}^{q+1}(u,v):=(-i)^{n-1}\left(\sum_{j=1}^{+\infty}d_{j}\nu_{j}^{q+1}\omega_{j}^{n-q}I^{j}_{T}(u_{q+1},v_{n-q})+\sum_{j=1}^{+\infty}\tilde{d_{j}}\nu_{j}^{n-q}\omega_{j}^{q+1}I^{j}_{T}(v_{n-q},u_{q+1})\right).$
The following lemma is strongly inspired by [18, Proposition $5.1$].
@@
$A=-i\sum_{j=1}^{+\infty}c_{j}\nu_{j}^{l+2}\omega_{j}^{l+1}I^{j}_{T}(u_{l+2},u_{l+1}),\hskip 14.22636ptB:=-\sum_{j=1}^{+\infty}c_{j}\nu_{j}^{l+1}\omega_{j}^{l+1}\int_{0}^{T}u_{l+2}(t)u_{l+1}(t)e^{i\omega_{K}t}\mathrm{d}t.$
@@
$A=-i\sum_{j=1}^{+\infty}c_{j}\nu_{j}^{l+2}\omega_{j}^{l+1}\int_{0}^{T}u_{l+2}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+(-1)^{l}R_{T,1}^{l+2}(u).$ (4.4)
$B=\frac{i\omega_{K}}{2}\sum_{j=1}^{+\infty}c_{j}\nu_{j}^{l+1}\omega_{j}^{l+1}\int_{0}^{T}u_{l+2}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t.$ (4.5)
@@
Now, we can follow the same approach with the crossed terms. More precisely,
@@
We deal with the case where $n$ is odd. Thanks to (3.11) and (3.13), we obtain: $\mathcal{G}_{T}(u,v)=-(A+B)$ with
@@
$A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)+(-1)^{l+1}\sum_{j=1}^{+\infty}d_{j}\nu_{j}^{l+1}\omega_{j}^{l+1}I^{j}_{T}(u_{l+1},v_{l+1})+\\ \sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\nu_{j}^{k}\omega_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\nu_{j}^{k+1}\omega_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k+1}).$ (4.8)
@@
$A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)-\sum_{j=1}^{+\infty}d_{j}\nu_{j}\omega_{j}I^{j}_{T}(u_{1},v_{1})+\\ -\sum_{j=1}^{+\infty}d_{j}\mathcal{I}_{T}(u_{1},v)-i\sum_{j=1}^{+\infty}d_{j}\nu_{j}\mathcal{I}_{T}(u_{1},v_{1}).$

### v4

@@ 1.6 Small-time continuously approximately reachable vector
Finally, let $(\mathscr{F}_{T})_{T>0}$ be a family of functions from $X\times E_{T}$ to $X$ for $T>0$. The goal is to prove that the map $\mathscr{F}_{T}$ in locally onto. To do this, we use the following definition; this is an equivalent of the notion of tangent vector introduced by Kawski in [36], but in infinite-dimensional case.
@@
A vector $\xi\in X$ is called a small-time continuously approximately reachable vector if there exists a continuous map $\Xi:[0,+\infty[\to X$ with $\Xi(0)=\xi$ such that for all $T>0$, there exist $C,\rho,s>0$ and a continuous map $z\in(-\rho,\rho)\mapsto u_{z}\in E_{T}$ such that
@@ 1.7 Plan of the paper
The paper is organized in the following way: in section $2$, we present a proof of a theorem of STLC thanks to quadratic terms for control-affine system in the finite-dimensional case, as a toy-model for the bilinear Schrödinger equation. In section $3$, we present classical properties about the bilinear Schrödinger equation. Finally, in section $4$, we give the proof of the main theorem of this article. Some elements of proof are developed in the appendix.
@@ 2.1 Framework and notations
For each $0<T_{1}<T$, $u,v\in L^{1}((T_{1},T),\mathbb{R})$, there exists a unique maximal mild solution to (2.1) with initial data $p\in\mathbb{R}^{d}$ at time $T_{1}$, which we will denote by $x(\cdot;(u,v),p)$. We will sometimes note $x(\cdot;(u,v),p,T_{1})$. As we are interested in small-time, and small controls, the solution is well-defined up to time $T$. The following concepts were introduced by Beauchard and Marbach in [10], in 2018.
@@
Let $f_{0},f_{1},f_{2}:\Omega\to\mathbb{R}^{d}$ be smooth vector fields on an open subset $\Omega$ of $\mathbb{R}^{d}$ and $f=\{f_{0},f_{1},f_{2}\}$.
@@ 2.3 A quadratic sufficient condition for STLC
Let $d,L\in\mathbb{N}^{*}$, $f_{0},f_{1},f_{2}$ be analytic vector fields on a neighborhood of 00 in $\mathbb{R}^{d}$, such that $f_{0}(0)=0$. Assume that $r:=\dim\left(S_{1}(f)(0)\right)<d$ and that there exist brackets $b_{r+1},\cdots,b_{d}\in\mathcal{B}_{2,good}$ such that $\underset{i\in\llbracket r+1,d\rrbracket}{\max}|b_{i}|=L$ and
@@ 2.4 Some tools for the proof
Let $\mathcal{B}\subset\operatorname{Br}(X)$ be a Hall set. The coordinates of the second kind associated with $\mathcal{B}$ is the unique family $(\xi_{b})_{b\in\mathcal{B}}$ of functionals +×$\mathbb{R}_{+}\times L^{1}_{\mathrm{loc}}(\mathbb{R}_{+},\mathbb{R})^{2}\to\mathbb{R}$ defined by induction as: for every $t>0$ and $u,v\in L^{1}((0,t),\mathbb{R})$,
  • $\xi_{X_{0}}(t,(u,v)):=t$,  $\xi_{X_{1}}(t,(u,v)):=u_{1}(t)$  and  $\xi_{X_{2}}(t,(u,v)):=v_{1}(t)$,
@@
For every $M\in\mathbb{N}^{*}$, there exists $C_{M}>0$ such that, for every $T>0$, $u,v\in L^{1}((0,T),\mathbb{R})$, $b\in\mathcal{B}$ with $n(b)\leqslant M$ and $t\in[0,T]$,
@@ 2.5 A quadratic moment problem
One may assume that $B=\left\{M_{\nu}^{1},\ M_{\nu}^{2},\ C_{\mu,\nu},\ (\mu,\nu)\in\llbracket 0,\mu^{*}\rrbracket\times\llbracket 0,\nu^{*}\rrbracket\right\}.$ One considers $N=\mu^{*}+\nu^{*}+1$. We are looking for $u$ and $v$ in the form of $u=\phi^{(N)}$ and $v=\psi^{(N)}$, with $\phi,\psi\in\mathcal{C}^{\infty}_{c}((0,1),\mathbb{R})$. One recalls that for $b\in\mathcal{B}_{1}$, $\eta_{b}=\xi_{b}$. Then, for every $\mu\in\llbracket 0,\mu^{*}\rrbracket$, $\eta_{M_{\mu}^{1}}(1,(u,v))=u_{\mu+1}(1)=0$, and $\eta_{M_{\mu}^{2}}(1,(u,v))=v_{\mu+1}(1)=0$, using (2.10). Thus, (2.18) is already verified for $b\in B\cap\mathcal{B}_{1}$.
Thanks to the choice of $N$, the equation (2.12) and (2.14) give: for all $(\mu,\nu)\in\llbracket 0,\mu^{*}\rrbracket\times\llbracket 0,\nu^{*}\rrbracket$,
$\eta_{C_{\mu,\nu}}(1,(u,v))=\xi_{C_{\mu,\nu}}(1,(u,v))=\int_{0}^{1}\frac{(1-t)^{\nu}}{\nu!}u_{\lfloor\frac{\mu}{2}\rfloor+1}(t)v_{\lfloor\frac{\mu+1}{2}\rfloor}(t)\mathrm{d}t.$
@@
Step 2: Construction of $u$ and $v$. One chooses $u:=\phi^{(N)}$, with the function $\phi$ obtained at step $1$. By construction of $\phi$, there exists $\psi\in\mathcal{C}^{\infty}_{c}((0,1),\mathbb{R})$, solution to the following moment problems
@@ 2.6 Tangent vector
Let $m\geqslant 0$, we prove that the system is $W^{m,\infty}_{0}-$STLC. Taking into account the assumption (2.7), we can consider $\mathbb{P}$, the linear projection on $\text{Span}(f_{b_{i}}(0))_{r+1\leqslant i\leqslant d}$ parallel to ${S}_{1}(f)(0)$. The purpose of this section is to prove the following proposition.
@@
Consequently, the definition of $(\bar{u},\bar{v})$ and (2.24) gives: for all $z\in(-\rho,\rho)$, for all $i\in\{1,2\}$, for all $k\in\llbracket 0,q_{j}\rrbracket$, $\eta_{M_{k}^{i}}(T_{1},(u_{z},v_{z}))=0$. Using (2.17) and (2.21) with $r=1$,
@@
Using the size of the controls given by (2.21) with $(k,r)=(0,1)$ and (2.25), one has
@@
Let $j\in\llbracket r+1,d\rrbracket$. For all $10<T_{1}<T<T_{1}+$ 1, there exist $C,\rho>0$ and a continuous map $z\mapsto(U_{z},V_{z})\in W^{m,\infty}_{0}(0,T)^{2}$ such that
@@
Note that $\mathbb{P}_{{S}_{1}(f)(0)}=\operatorname{Id}-\mathbb{P}$. Then,
@@
$e^{(T-T_{1})H_{0}}\left(\operatorname{Id}-\mathbb{P}\right)\left(x(T_{1};(u_{z},v_{z}),0,0)\right)\in{S}_{1}(f)(0)=\ker(\mathbb{P}),$
@@
$\begin{gathered}\left\|\mathbb{P}\left(\mathcal{Z}_{2}(T;f,(\tilde{u}_{z},\tilde{v}_{z}))\right)\right\|_{\infty}=\mathcal{O}\left(\left\|(\tilde{u}_{z},\tilde{v}_{z})\right\|_{L^{1}}^{2}\right)=\mathcal{O}\left(\left\|(\tilde{u}_{z},\tilde{v}_{z})\right\|_{W^{m,\infty}}^{2}\right)=\mathcal{O}\left(|z|^{2\delta_{j}}\right),\\ \left\|\mathcal{Z}_{2}(T;f,(\tilde{u}_{z},\tilde{v}_{z}))\right\|_{\infty},\left\|D\mathcal{Z}_{2}(T;f,(\tilde{u}_{z},\tilde{v}_{z}))\right\|_{\infty}=\mathcal{O}\left(\left\|(\tilde{u}_{z},\tilde{v}_{z})\right\|_{W^{m,\infty}}\right)=\mathcal{O}\left(|z|^{\delta_{j}}\right).\end{gathered}$
Therefore, thanks to the estimates (2.38), one has
@@
using (2.32) and $1+s_{j}\leqslant 2\delta_{j}$. To conclude, the equations (2.36), (2.37) and (2.39) lead to the wondered inequality (2.30). ∎
@@ 5. $(A_{5})$
    Finally, Proposition 2.24 shows that, for all $j\in\llbracket r+1,d\rrbracket$, $f_{b_{j}}(0)$ is a small-time $W^{m,\infty}_{0}$-continuously approximately reachable vector. Then, as a supplementary space of $H$ is given by $\text{Span}\left(\left(f_{b_{j}}(0)\right)_{r+1\leqslant j\leqslant d}\right)$, this condition is verified.
@@ 3 Preliminaries on the Schrödinger equation
In all the document, we will note $\omega_{j}:=\lambda_{j}-\lambda_{1}$, and $\nu_{j}:=\lambda_{K}-\lambda_{j}$, with $j\geqslant 1$, $K\geqslant 2$. For $\varphi\in L^{2}(0,1)$, $t\in\mathbb{R}$, one defines
@@ 3.1 Well-posedness
Let $\mu$ be such a function. For all $\varphi\in H^{2k}_{(0)}(0,1)$, for all $l\in\{0,\cdots,k-1\}$, one has, thanks to Leibniz formula
@@
$0=\left(\mu\varphi_{\ell}\right)^{(2k-2)}(0)=\sum_{j=0}^{k-2}\binom{2k-2}{2j+1}\mu^{(2j+1)}(0)\delta_{j,\ell}=\binom{2k-2}{2\ell+1}\mu^{(2\ell+1)}(0).$
Using the same strategy with $x=1$, we obtain the result. ∎
@@ 3.2 Expansion of the solution
$\left\{\begin{array}[]{lr}i\partial_{t}\xi=-\partial_{x}^{2}\xi-(u\mu_{1}+v\mu_{2})\Psi,&(0,T)\times(0,1),\\ \xi(\cdot,0)=\xi(\cdot,1)=0,&(0,T),\\ \xi(0,\cdot)=0,&(0,1).\end{array}\right.$ (3.10)
@@
Third step: cubic remainder. Once again, one defines $\Lambda:t\in[0,T]\mapsto\psi(t;(u,v),\varphi_{1})-\psi_{1}(t)-\Psi(t)-\xi(t)$. The function $\Lambda$ is solution to the following PDE
$\left\{\begin{array}[]{lr}i\partial_{t}\Lambda=-\partial_{x}^{2}\Lambda-(u\mu_{1}+v\mu_{2})\left(\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right),&(0,T)\times(0,1),\\ \Lambda(\cdot,0)=\Lambda(\cdot,1)=0,&(0,T),\\ \Lambda(0,\cdot)=0,&(0,1).\end{array}\right.$
One uses the estimate (3.3) with $m=0$ in this equation to obtain
$\left\|\Lambda\right\|_{L^{\infty}\left((0,T),H^{2p+3}_{(0)}\right)}\leqslant C\left\|(u\mu_{1}+v\mu_{2})\left(\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right)\right\|_{L^{2}((0,T),H^{2p+3})}.$
Using the algebra structure of the Sobolev space $H^{2p+3}(0,1)$,
$\left\|u\mu_{1}\left(\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right)\right\|_{L^{2}\left((0,T),H^{2p+3}\right)}\\ \leqslant C\left\|\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right\|_{L^{\infty}\left((0,T),H^{2p+3}_{(0)}\right)}\left\|\mu_{1}\right\|_{H^{2p+3}}\left\|u\right\|_{L^{2}(0,T)}.$ (3.18)
Then, using the inequality proved in the second step in (3.18), we obtain
$\left\|\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi-\xi\right\|_{L^{\infty}\left((0,T),H^{2p+3}_{(0)}\right)}=\mathcal{O}\left(\left\|u\right\|_{L^{2}(0,T)}^{3}+\left\|v\right\|_{L^{2}(0,T)}^{3}\right).$
@@ 3.3 Control in projection
Let $p,m,k\geqslant 0$, $K\geqslant 2$, such that $k\leqslant p$, $\mu_{1},\mu_{2}$ functions satisfying $\mathbf{(H)_{reg}}$, and $\mathbf{(H)_{lin,K,2}}$. Then, the bilinear Schrödinger equation (1.1) is $H^{m}_{0}-$STLC in projection on $\mathcal{H}$ around the ground state in $H^{2(p+m)+3}_{(0)}$. More precisely, for any $T_{1}<T$, there exist $\delta,C>0$, and a $\mathcal{C}^{1}$ map $\Gamma_{T_{1},T}:\mathcal{V}_{T_{1}}\times\mathcal{V}_{T}\to H^{m}_{0}((T_{1},T),\mathbb{R})^{2}$, where
@@ 4 Proof of the main theorem
The proof of the main theorem is divided in the following steps.
@@
In order to extract the leading terms of the dynamic of the system (as in the finite-dimensional case, with the Magnus-type representation formula), we manipulate the expression given by (3.14). This is the purpose of the following subsection.
@@ 4.1 Asymptotic estimates on the quadratic term of the solution
We define $q:=\lfloor\frac{n}{2}\rfloor$. This notation will be used throughout the rest of the article. As a remainder, the sequence $(c_{j})_{j\geqslant 1}$ is defined in Section 1.3 as $c_{j}=\left\langle\mu_{1}\varphi_{1},\varphi_{j}\right\rangle\left\langle\mu_{1}\varphi_{j},\varphi_{K}\right\rangle$. The goal of this subsection is to prove the following proposition.
@@
Let $u,v\in L^{2}((0,T),\mathbb{R})$ be such that $u_{i}(T)=v_{i}(T)=0$ for $11\leqslant i\leqslant q+$ 1. Furthermore, assume that $\mathbf{(H)_{reg}}$, $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$ and $\mathbf{(H)_{quad,K,3}}$ hold. Then,
@@
$R_{T,1}^{q+1}(u):=\displaystyle(-1)^{q}\sum_{j=1}^{+\infty}c_{j}\omega_{j}^{q+1}\nu_{j}^{q+1}I^{j}_{T}(u_{q+1},u_{q+1}),$
$R_{T,2}^{q+1}(v):=\displaystyle(-1)^{q}\sum_{j=1}^{+\infty}\tilde{c_{j}}\omega_{j}^{q+1}\nu_{j}^{q+1}I^{j}_{T}(v_{q+1},v_{q+1}),$
$\rho_{T}^{q+1}(u,v):=(-i)^{n-1}\left(\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{n-q}\nu_{j}^{q+1}I^{j}_{T}(u_{q+1},v_{n-q})+\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{q+1}\nu_{j}^{n-q}I^{j}_{T}(v_{n-q},u_{q+1})\right).$
The proof is divided into several lemmas. The first one, strongly inspired by [18, Proposition $5.1$], focuses on the first term on the right-hand side of the decomposition given in (3.14).
@@
$A=-i\sum_{j=1}^{+\infty}c_{j}\omega_{j}^{l+1}\nu_{j}^{l+2}I^{j}_{T}(u_{l+2},u_{l+1}),\hskip 28.45274ptB:=-\sum_{j=1}^{+\infty}c_{j}\omega_{j}^{l+1}\nu_{j}^{l+1}\mathcal{I}_{T}(u_{l+2},u_{l+1}).$
@@
$A=-i\sum_{j=1}^{+\infty}c_{j}\omega_{j}^{l+1}\nu_{j}^{l+2}\int_{0}^{T}u_{l+2}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+(-1)^{l}R_{T,1}^{l+2}(u).$ (4.4)
$B=\frac{i\omega_{K}}{2}\sum_{j=1}^{+\infty}c_{j}\omega_{j}^{l+1}\nu_{j}^{l+1}\int_{0}^{T}u_{l+2}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t.$ (4.5)
@@
Now, we can follow the same approach with the crossed terms. More precisely, we focus on the second term on the right-hand side of the decomposition given by (3.14).
@@
We deal with the case where $n$ is odd. Thanks to (3.11) and (3.13), we obtain $\mathcal{G}_{T}(u,v)=-(A+B)$ with
@@
$A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)+(-1)^{l+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{l+1}\nu_{j}^{l+1}I^{j}_{T}(u_{l+1},v_{l+1})+\\ \sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{k}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{k}\nu_{j}^{k+1}\mathcal{I}_{T}(u_{k+1},v_{k+1}).$ (4.8)
@@
$A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)-\sum_{j=1}^{+\infty}d_{j}\omega_{j}\nu_{j}I^{j}_{T}(u_{1},v_{1})+\\ -\sum_{j=1}^{+\infty}d_{j}\mathcal{I}_{T}(u_{1},v)-i\sum_{j=1}^{+\infty}d_{j}\nu_{j}\mathcal{I}_{T}(u_{1},v_{1}).$